# Yang-Lee Quantum Criticality in Various Dimensions

### v3

@@ Appendix A Characters of $W(F_{4})$
$\textbf{1}(\phi_{1,0})$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
$\textbf{1}^{\prime}(\phi_{1,12}^{\prime\prime})$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $-1$ $-1$ $-1$ $-1$
$\textbf{1}^{\prime\prime}(\phi_{1,12}^{\prime})$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $-1$ $-1$ $-1$ $-1$
$\textbf{1}^{\prime\prime\prime}(\phi_{1,24})$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
$\textbf{2}(\phi_{2,4}^{\prime\prime})$ $2$ $2$ $2$ $2$ $2$ $2$ $-1$ $-1$ $-1$ $-1$ $-1$ 00 00 00 00
$\textbf{2}^{\prime}(\phi_{2,4}^{\prime})$ $2$ $2$ $2$ $-1$ $-1$ $2$ $2$ $2$ $-1$ $-1$ $-1$ 00 00 00 00
$\textbf{2}^{\prime\prime}(\phi_{2,16}^{\prime})$ $2$ $2$ $2$ $2$ $2$ $2$ $-1$ $-1$ $-1$ $-1$ $-1$ 00 00 00 00
$\textbf{2}^{\prime\prime\prime}(\phi_{2,16}^{\prime\prime})$ $2$ $2$ $2$ $-1$ $-1$ $2$ $2$ $2$ $-1$ $-1$ $-1$ 00 00 00 00
$\textbf{4}(\phi_{4,13})$ $4$ $-4$ 00 $1$ $-1$ 00 $1$ $-1$ $-2$ $2$ 00 00 $2$ $-2$ 00
$\textbf{4}^{\prime}(\phi_{4,8})$ $4$ $4$ $4$ $-2$ $-2$ $4$ $-2$ $-2$ $1$ $1$ $1$ 00 00 00 00
$\textbf{4}^{\prime\prime}(\phi_{4,1})$ $4$ $-4$ 00 $1$ $-1$ 00 $1$ $-1$ $-2$ $2$ 00 00 $2$ $-2$ 00
$\textbf{4}^{\prime\prime\prime}(\phi_{4,7}^{\prime\prime})$ $4$ $-4$ 00 $1$ $-1$ 00 $1$ $-1$ $-2$ $2$ 00 00 $-2$ $2$ 00
$\textbf{4}^{\prime\prime\prime\prime}(\phi_{4,7}^{\prime})$ $4$ $-4$ 00 $1$ $-1$ 00 $1$ $-1$ $-2$ $2$ 00 00 $-2$ $2$ 00
$\textbf{6}(\phi_{6,6}^{\prime})$ $6$ $6$ $-2$ 00 00 $2$ 00 00 $3$ $3$ $-1$ $2$ $-2$ $-2$ 00
$\textbf{6}^{\prime}(\phi_{6,6}^{\prime\prime})$ $6$ $6$ $-2$ 00 00 $2$ 00 00 $3$ $3$ $-1$ $-2$ $2$ $2$ 00
$\textbf{8}(\phi_{8,9}^{\prime})$ $8$ $-8$ 00 $2$ $-2$ 00 $-1$ $1$ $2$ $-2$ 00 00 00 00 00
$\textbf{8}^{\prime}(\phi_{8,9}^{\prime\prime})$ $8$ $-8$ 00 $-1$ $1$ 00 $2$ $-2$ $2$ $-2$ 00 00 00 00 00
$\textbf{8}^{\prime\prime}(\phi_{8,3}^{\prime\prime})$ $8$ $-8$ 00 $2$ $-2$ 00 $-1$ $1$ $2$ $-2$ 00 00 00 00 00
$\textbf{8}^{\prime\prime\prime}(\phi_{8,3}^{\prime})$ $8$ $-8$ 00 $-1$ $1$ 00 $2$ $-2$ $2$ $-2$ 00 00 00 00 00
$\textbf{9}(\phi_{9,2})$ $9$ $9$ $1$ 00 00 $-3$ 00 00 00 00 00 $1$ $1$ $1$ $-1$
$\textbf{9}^{\prime}(\phi_{9,6}^{\prime\prime})$ $9$ $9$ $1$ 00 00 $-3$ 00 00 00 00 00 $-1$ $-1$ $-1$ $1$
$\textbf{9}^{\prime\prime}(\phi_{9,6}^{\prime})$ $9$ $9$ $1$ 00 00 $-3$ 00 00 00 00 00 $-1$ $-1$ $-1$ $1$
$\textbf{9}^{\prime\prime\prime}(\phi_{9,10})$ $9$ $9$ $1$ 00 00 $-3$ 00 00 00 00 00 $1$ $1$ $1$ $-1$
$\textbf{12}(\phi_{12,4})$ $12$ $12$ $-4$ 00 00 $4$ 00 00 $-3$ $-3$ $1$ 00 00 00 00
$\textbf{16}(\phi_{16,1})$ $16$ $-16$ 00 $-2$ $2$ 00 $-2$ $2$ $-2$ $2$ 00 00 00 00 00
@@
$\textbf{1}(\phi_{1,0})$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$ $1$
$\textbf{1}^{\prime}(\phi_{1,12}^{\prime\prime})$ $1$ $1$ $1$ $1$ $1$ $-1$ $-1$ $-1$ $-1$ $-1$
$\textbf{1}^{\prime\prime}(\phi_{1,12}^{\prime})$ $-1$ $-1$ $-1$ $-1$ $-1$ $1$ $1$ $1$ $1$ $1$
$\textbf{1}^{\prime\prime\prime}(\phi_{1,24})$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$
$\textbf{2}(\phi_{2,4}^{\prime\prime})$ $2$ $2$ $-1$ $-1$ $2$ 00 00 00 00 00
$\textbf{2}^{\prime}(\phi_{2,4}^{\prime})$ 00 00 00 00 00 $2$ $2$ $-1$ $-1$ $2$
$\textbf{2}^{\prime\prime}(\phi_{2,16}^{\prime})$ $-2$ $-2$ $1$ $1$ $-2$ 00 00 00 00 00
$\textbf{2}^{\prime\prime\prime}(\phi_{2,16}^{\prime\prime})$ 00 00 00 00 00 $-2$ $-2$ $1$ $1$ $-2$
$\textbf{4}(\phi_{4,13})$ $-2$ $2$ $1$ $-1$ 00 $-2$ $2$ $1$ $-1$ 00
$\textbf{4}^{\prime}(\phi_{4,8})$ 00 00 00 00 00 00 00 00 00 00
$\textbf{4}^{\prime\prime}(\phi_{4,1})$ $2$ $-2$ $-1$ $1$ 00 $2$ $-2$ $-1$ $1$ 00
$\textbf{4}^{\prime\prime\prime}(\phi_{4,7}^{\prime\prime})$ $2$ $-2$ $-1$ $1$ 00 $-2$ $2$ $1$ $-1$ 00
$\textbf{4}^{\prime\prime\prime\prime}(\phi_{4,7}^{\prime})$ $-2$ $2$ $1$ $-1$ 00 $2$ $-2$ $-1$ $1$ 00
$\textbf{6}(\phi_{6,6}^{\prime})$ 00 00 00 00 00 00 00 00 00 00
$\textbf{6}^{\prime}(\phi_{6,6}^{\prime\prime})$ 00 00 00 00 00 00 00 00 00 00
$\textbf{8}(\phi_{8,9}^{\prime})$ $-4$ $4$ $-1$ $1$ 00 00 00 00 00 00
$\textbf{8}^{\prime}(\phi_{8,9}^{\prime\prime})$ 00 00 00 00 00 $-4$ $4$ $-1$ $1$ 00
$\textbf{8}^{\prime\prime}(\phi_{8,3}^{\prime\prime})$ $4$ $-4$ $1$ $-1$ 00 00 00 00 00 00
$\textbf{8}^{\prime\prime\prime}(\phi_{8,3}^{\prime})$ 00 00 00 00 00 $4$ $-4$ $1$ $-1$ 00
$\textbf{9}(\phi_{9,2})$ $3$ $3$ 00 00 $-1$ $3$ $3$ 00 00 $-1$
$\textbf{9}^{\prime}(\phi_{9,6}^{\prime\prime})$ $3$ $3$ 00 00 $-1$ $-3$ $-3$ 00 00 $1$
$\textbf{9}^{\prime\prime}(\phi_{9,6}^{\prime})$ $-3$ $-3$ 00 00 $1$ $3$ $3$ 00 00 $-1$
$\textbf{9}^{\prime\prime\prime}(\phi_{9,10})$ $-3$ $-3$ 00 00 $1$ $-3$ $-3$ 00 00 $1$
$\textbf{12}(\phi_{12,4})$ 00 00 00 00 00 00 00 00 00 00
$\textbf{16}(\phi_{16,1})$ 00 00 00 00 00 00 00 00 00 00
@@ Appendix B Constructing Spinning Correlation Functions in Fuzzy Sphere
where $I_{\mu\nu}=\delta_{\mu\nu}-2\frac{x_{\mu}x_{\nu}}{x^{2}}$. Strictly speaking, the ket-state is not quite the conjugation but the left eigenvector and the inversion will also take ${\cal O}\mapsto{\cal O}^{\dagger}$. In YL model the left and right eigenvector are related by complex conjugation which cancels the conjugation of the operator if it is not real. The net effect is that we do not need to worry about the conjugation of ${\cal O}$. By extrapolating ${\cal O}_{1}$ and ${\cal O}_{3}$ to $\infty$ and 00, we get a much cleaner expression

### v4

@@ Appendix A Characters of $W(F_{4})$
$\textbf{1}(\phi_{1,0})$ 11 11 11 11 11 11 11 11 11 11 11 11 11 11 11
$\textbf{1}^{\prime}(\phi_{1,12}^{\prime\prime})$ 11 11 11 11 11 11 11 11 11 11 11 1-1 1-1 1-1 1-1
$\textbf{1}^{\prime\prime}(\phi_{1,12}^{\prime})$ 11 11 11 11 11 11 11 11 11 11 11 1-1 1-1 1-1 1-1
$\textbf{1}^{\prime\prime\prime}(\phi_{1,24})$ 11 11 11 11 11 11 11 11 11 11 11 11 11 11 11
$\textbf{2}(\phi_{2,4}^{\prime\prime})$ 22 22 22 22 22 22 1-1 1-1 1-1 1-1 1-1 0 0 0 0
$\textbf{2}^{\prime}(\phi_{2,4}^{\prime})$ 22 22 22 1-1 1-1 22 22 22 1-1 1-1 1-1 0 0 0 0
$\textbf{2}^{\prime\prime}(\phi_{2,16}^{\prime})$ 22 22 22 22 22 22 1-1 1-1 1-1 1-1 1-1 0 0 0 0
$\textbf{2}^{\prime\prime\prime}(\phi_{2,16}^{\prime\prime})$ 22 22 22 1-1 1-1 22 22 22 1-1 1-1 1-1 0 0 0 0
$\textbf{4}(\phi_{4,13})$ 44 4-4 0 11 1-1 0 11 1-1 2-2 22 0 0 22 2-2 0
$\textbf{4}^{\prime}(\phi_{4,8})$ 44 44 44 2-2 2-2 44 2-2 2-2 11 11 11 0 0 0 0
$\textbf{4}^{\prime\prime}(\phi_{4,1})$ 44 4-4 0 11 1-1 0 11 1-1 2-2 22 0 0 22 2-2 0
$\textbf{4}^{\prime\prime\prime}(\phi_{4,7}^{\prime\prime})$ 44 4-4 0 11 1-1 0 11 1-1 2-2 22 0 0 2-2 22 0
$\textbf{4}^{\prime\prime\prime\prime}(\phi_{4,7}^{\prime})$ 44 4-4 0 11 1-1 0 11 1-1 2-2 22 0 0 2-2 22 0
$\textbf{6}(\phi_{6,6}^{\prime})$ 66 66 2-2 0 0 22 0 0 33 33 1-1 22 2-2 2-2 0
$\textbf{6}^{\prime}(\phi_{6,6}^{\prime\prime})$ 66 66 2-2 0 0 22 0 0 33 33 1-1 2-2 22 22 0
$\textbf{8}(\phi_{8,9}^{\prime})$ 88 8-8 0 22 2-2 0 1-1 11 22 2-2 0 0 0 0 0
$\textbf{8}^{\prime}(\phi_{8,9}^{\prime\prime})$ 88 8-8 0 1-1 11 0 22 2-2 22 2-2 0 0 0 0 0
$\textbf{8}^{\prime\prime}(\phi_{8,3}^{\prime\prime})$ 88 8-8 0 22 2-2 0 1-1 11 22 2-2 0 0 0 0 0
$\textbf{8}^{\prime\prime\prime}(\phi_{8,3}^{\prime})$ 88 8-8 0 1-1 11 0 22 2-2 22 2-2 0 0 0 0 0
$\textbf{9}(\phi_{9,2})$ 99 99 11 0 0 3-3 0 0 0 0 0 11 11 11 1-1
$\textbf{9}^{\prime}(\phi_{9,6}^{\prime\prime})$ 99 99 11 0 0 3-3 0 0 0 0 0 1-1 1-1 1-1 11
$\textbf{9}^{\prime\prime}(\phi_{9,6}^{\prime})$ 99 99 11 0 0 3-3 0 0 0 0 0 1-1 1-1 1-1 11
$\textbf{9}^{\prime\prime\prime}(\phi_{9,10})$ 99 99 11 0 0 3-3 0 0 0 0 0 11 11 11 1-1
$\textbf{12}(\phi_{12,4})$ $12$ $12$ 4-4 0 0 44 0 0 3-3 3-3 11 0 0 0 0
$\textbf{16}(\phi_{16,1})$ $16$ 16-16 0 2-2 22 0 2-2 22 2-2 22 0 0 0 0 0
@@
$\textbf{1}(\phi_{1,0})$ 11 11 11 11 11 11 11 11 11 11
$\textbf{1}^{\prime}(\phi_{1,12}^{\prime\prime})$ 11 11 11 11 11 1-1 1-1 1-1 1-1 1-1
$\textbf{1}^{\prime\prime}(\phi_{1,12}^{\prime})$ 1-1 1-1 1-1 1-1 1-1 11 11 11 11 11
$\textbf{1}^{\prime\prime\prime}(\phi_{1,24})$ 1-1 1-1 1-1 1-1 1-1 1-1 1-1 1-1 1-1 1-1
$\textbf{2}(\phi_{2,4}^{\prime\prime})$ 22 22 1-1 1-1 22 0 0 0 0 0
$\textbf{2}^{\prime}(\phi_{2,4}^{\prime})$ 0 0 0 0 0 22 22 1-1 1-1 22
$\textbf{2}^{\prime\prime}(\phi_{2,16}^{\prime})$ 2-2 2-2 11 11 2-2 0 0 0 0 0
$\textbf{2}^{\prime\prime\prime}(\phi_{2,16}^{\prime\prime})$ 0 0 0 0 0 2-2 2-2 11 11 2-2
$\textbf{4}(\phi_{4,13})$ 2-2 22 11 1-1 0 2-2 22 11 1-1 0
$\textbf{4}^{\prime}(\phi_{4,8})$ 0 0 0 0 0 0 0 0 0 0
$\textbf{4}^{\prime\prime}(\phi_{4,1})$ 22 2-2 1-1 11 0 22 2-2 1-1 11 0
$\textbf{4}^{\prime\prime\prime}(\phi_{4,7}^{\prime\prime})$ 22 2-2 1-1 11 0 2-2 22 11 1-1 0
$\textbf{4}^{\prime\prime\prime\prime}(\phi_{4,7}^{\prime})$ 2-2 22 11 1-1 0 22 2-2 1-1 11 0
$\textbf{6}(\phi_{6,6}^{\prime})$ 0 0 0 0 0 0 0 0 0 0
$\textbf{6}^{\prime}(\phi_{6,6}^{\prime\prime})$ 0 0 0 0 0 0 0 0 0 0
$\textbf{8}(\phi_{8,9}^{\prime})$ 4-4 44 1-1 11 0 0 0 0 0 0
$\textbf{8}^{\prime}(\phi_{8,9}^{\prime\prime})$ 0 0 0 0 0 4-4 44 1-1 11 0
$\textbf{8}^{\prime\prime}(\phi_{8,3}^{\prime\prime})$ 44 4-4 11 1-1 0 0 0 0 0 0
$\textbf{8}^{\prime\prime\prime}(\phi_{8,3}^{\prime})$ 0 0 0 0 0 44 4-4 11 1-1 0
$\textbf{9}(\phi_{9,2})$ 33 33 0 0 1-1 33 33 0 0 1-1
$\textbf{9}^{\prime}(\phi_{9,6}^{\prime\prime})$ 33 33 0 0 1-1 3-3 3-3 0 0 11
$\textbf{9}^{\prime\prime}(\phi_{9,6}^{\prime})$ 3-3 3-3 0 0 11 33 33 0 0 1-1
$\textbf{9}^{\prime\prime\prime}(\phi_{9,10})$ 3-3 3-3 0 0 11 3-3 3-3 0 0 11
$\textbf{12}(\phi_{12,4})$ 0 0 0 0 0 0 0 0 0 0
$\textbf{16}(\phi_{16,1})$ 0 0 0 0 0 0 0 0 0 0
@@ Appendix B Constructing Spinning Correlation Functions in Fuzzy Sphere
where $I_{\mu\nu}=\delta_{\mu\nu}-2\frac{x_{\mu}x_{\nu}}{x^{2}}$. Strictly speaking, the ket-state is not quite the conjugation but the left eigenvector and the inversion will also take ${\cal O}\mapsto{\cal O}^{\dagger}$. In YL model the left and right eigenvector are related by complex conjugation which cancels the conjugation of the operator if it is not real. The net effect is that we do not need to worry about the conjugation of $\cal O$}. By extrapolating ${\cal O}_{1}$ and ${\cal O}_{3}$ to $\infty$ and 0, we get a much cleaner expression